write the equation in standard form with integer coefficients -8y=20+2/5x
Answer by haileytucki(390)

-8y=20+(2)/(5)*x
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
20+(2)/(5)*x=-8y
Multiply (2)/(5) by x to get (2x)/(5).
20+(2x)/(5)=-8y
Since 20 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 20 from both sides.
(2x)/(5)=-20-8y
Move all terms not containing x to the right-hand side of the equation.
(2x)/(5)=-8y-20
Multiply each term in the equation by 5.
(2x)/(5)*5=-8y*5-20*5
Simplify the left-hand side of the equation by canceling the common factors.
2x=-8y*5-20*5
Simplify the right-hand side of the equation by multiplying out all the terms.
2x=-40y-100
Divide each term in the equation by 2.
(2x)/(2)=-(40y)/(2)-(100)/(2)
Simplify the left-hand side of the equation by canceling the common factors.
x=-(40y)/(2)-(100)/(2)
Simplify the right-hand side of the equation by simplifying each term.
x=-20y-50
To rewrite as a function of y (e.g. f(y)), write the equation so that x is by itself on one side of the equal sign and an expression involving only y is on the other side.
f(y)=-20y-50